OpenStudy (raffle_snaffle):

7. Solar-heat installations successfully reduce the utility bill 60% of the time. What is the probability that at least 9 out of 10 solar-heat installations are successful and will reduce the utility bill? a. 0.0464 b. 0.9432 c. 0.0403 d. 0.8429

OpenStudy (raffle_snaffle):

@jim_thompson5910

OpenStudy (raffle_snaffle):

I used a binomial distr. on this problem

jimthompson5910 (jim_thompson5910):

yes you'll need to use a binomial cdf

OpenStudy (raffle_snaffle):

i did and I got 0.9987 but the solution key says it's a)

jimthompson5910 (jim_thompson5910):

how did you get 0.9987 ?

OpenStudy (raffle_snaffle):

oops wrong problem.. I mean 0.09473

OpenStudy (raffle_snaffle):

i use p = 0.60 and y = 9 and n = 10

jimthompson5910 (jim_thompson5910):

which calculator did you use?

OpenStudy (raffle_snaffle):

I used the 89 and yes I know about nCr(10,9)*p^y*q^(n-y)

OpenStudy (raffle_snaffle):

That is what I used..

OpenStudy (raffle_snaffle):

oh shoot I did it over again and got c.

jimthompson5910 (jim_thompson5910):

that's the probability of getting exactly 9 out of 10 you aren't considering 8 out of 10, 7 out of 10, etc and then adding them up

jimthompson5910 (jim_thompson5910):

sorry I meant to say just 9 out of 10 and 10 out of 10

OpenStudy (raffle_snaffle):

can I still use the nCr command?

jimthompson5910 (jim_thompson5910):

probability of getting exactly 9 out of 10 = 0.0403 probability of getting exactly 10 out of 10 = x the answer will be equal to `0.0403 + x`

OpenStudy (raffle_snaffle):

oh okay i see

OpenStudy (raffle_snaffle):

got it 0.04636

OpenStudy (raffle_snaffle):

so we are looking at the ones that failed?

jimthompson5910 (jim_thompson5910):

`so we are looking at the ones that failed? ` what do you mean?

jimthompson5910 (jim_thompson5910):

you should get 0.0463574016 which rounds to 0.04636

OpenStudy (raffle_snaffle):

I am sorry I got that but i rounded. nvm about my statement above it doesn't make sense. why do we look at 9 out of 10 and 10 out of 10?

jimthompson5910 (jim_thompson5910):

`at least 9 out of 10` = (exactly 9 out of 10) + (exactly 10 out of 10)

jimthompson5910 (jim_thompson5910):

saying "at least 9" means we have 9 or more so we have 9 or we have 10. We can't have any more than 10 because n = 10 is the sample size

OpenStudy (raffle_snaffle):

I see so P(Y>=9)

jimthompson5910 (jim_thompson5910):

correct

jimthompson5910 (jim_thompson5910):

\(\Large P(Y \ge 9)\) is the same as \(\Large P(Y = 9) + P(Y = 10)\)